Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below Indicate fie missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. overhead rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.) Case A Case C Direct materials used Direct labor Manufacturing overhead applied $88,000 $66,650 8,000 46,400 147,950 150,700 C0) 210,650 Work in process 1/1/17 Total cost of work in process Work in process 12/31/17 Cost of goods manufactured 18,300 18,700 204,800 13,400 193,500 224,800
Answer #1

Case A

Case B

Case C

Direct material Used

$      43,550.00

$      88,000.00

$      66,650.00

Direct labor

$      58,000.00

$ 1,50,700.00

$      80,000.00

Manufacturing overhead applied

$      46,400.00

$ 1,20,560.00

$      64,000.00

Total Manufacturing cost'

$ 1,47,950.00

$ 3,59,260.00

$ 2,10,650.00

Wouk in process 1/1/17

$      56,850.00

$      18,300.00

$      18,700.00

Total work in process

$ 2,04,800.00

$ 3,77,560.00

$ 2,29,350.00

Work in process 12/31/17

$      11,300.00

$      13,400.00

$        4,550.00

Cost of Goods manufactured

$ 1,93,500.00

$ 3,64,160.00

$ 2,24,800.00

Working

Case A

Case B

Case C

Direct material Used

=147950-58000-46400

88000

66650

Direct labor

58000

150700

=+(210650-66650)/180*100

Manufacturing overhead applied

46400

=150700*80%

=+(210650-66650)/180*80

Total Manufacturing cost'

147950

=88000+150700+120560

210650

Wouk in process 1/1/17

=204800-147950

18300

18700

Total work in process

204800

=359260+18300

=210650+18700

Work in process 12/31/17

=204800-193500

13400

=229350-224800

Cost of Goods manufactured

193500

=377560-13400

224800

Direct labor cost in case A

$      58,000.00

Manufacturing overhead applied in case A

$      46,400.00

Rate of Manufacturing overhead to direct labor

80%